OpenStudy (anonymous):

Evaluate the indefinite integral (a+bx^35)/(sqrt of (36ax+bx^36))dx u=36ax+bx^36 and (1/36)du=(a+bx^35)dx please help

hartnn (hartnn):

good so far! now plug all those into your original integral! and make a new integral

hartnn (hartnn):

numerator = (a+bx^35) dx = du/36 denominator = sqrt of (36ax+bx^36) = sqrt u

OpenStudy (anonymous):

(1/36)*(u^(1/2)/(1/2))+C

hartnn (hartnn):

before integration, its sqrt u ,right ???

hartnn (hartnn):

so, u^(1/2) before integration what will you get after integration ?

hartnn (hartnn):

cos ? 13/2 ? where does those come from ? :O

OpenStudy (anonymous):

wrong problem

hartnn (hartnn):